OpenStudy (anonymous):

How do I solve for x? (5)^2x=20

OpenStudy (tkhunny):

1) Write very clearly what you mean. \(5^{2}\cdot x = 20\) or \(5^{2x} = 20\)?

OpenStudy (anonymous):

My apologizes; I don't know how to represent equations over the laptop, but it is indeed the second one you typed.

OpenStudy (anonymous):

\[5^{2x}=20\]

OpenStudy (tkhunny):

No worries. Learning to communicate in this medium sometomes takes a little time. I see you have almost figured it all out. Good work. Have you considered logarithms?

OpenStudy (anonymous):

How do logarithms come into play with this equation? -I mean that is the subject we are currently in, but I'm struggling with how they function in general. I find all this log/ln/natural base e stuff to be quite a mess.

OpenStudy (tkhunny):

Logarithms ARE exponents. Logarithms are also useful for transforming expressions and saving relationships. If \(5^{2x} = 20\), then \(\log\left(5^{2x}\right) = log(20)\)

OpenStudy (anonymous):

I'm sorry, I don't know how to plug that into my calculator, is there a step process to solve that?

OpenStudy (anonymous):

1) In(25^x)=In(20) 2) x*In(25)=In(20) then, x=In(20)/In(25)=0.9307

OpenStudy (anonymous):

I got that same number (0.9307) as well. Didn't seem right to me though; apparently it is! Thanks so much both of you.